OpenStudy (anonymous):

Solve: 3(x + 1) + 6 = 33

OpenStudy (anonymous):

First you have to follow PEMDAS. Parenthesis, Exponents, Multiply, Divide, Add, Subtract. So, since there are parenthesis, we start with distribution to remove the parenthesis. 3(x + 1) 3x + 3 because 3*x = 3x and 3*1 = 3. 3x + 1 + 6 = 33. Add like factors. 1 + 6 = 7 3x + 7 = 33. Now like your last one, you need the variable on the opposite side, so we have to move the 7 to the other side. Once again, its a positive number so we subtract it, and add it to the other side. 3x + 7 - 7 = 33 + 7 3x = 40 And then you have to get the variable alone, and to do that you must do the opposite of the operator connecting the var. In this case, since 3x is multiplication, you must divide. x = 40/3

OpenStudy (aroub):

3x+3+6=33 3x+9=33 3x=24 x=8